If the perimeter of a rectangle is 24 3/5 and one side of the rectangle is 4 4/5 what are the other sides?
Answer by rfer(16322)

2x+294 4/5)=24 3/5
2x+2(24/5)=123/5
2x=123/5-48/5
2x=75/5
x=(75/5)/2
x=75/5*1/2
x=75/10
x=7 1/2